OpenStudy (anonymous):

Find tan θ if cos θ = one divided by four and sin θ < 0. negative square root of fifteen 4 negative square root of fifteen divided by fifteen negative square root of seventeen

OpenStudy (anonymous):

@geekfromthefutur @Griff85000 @mathmath333 any help on this one?

OpenStudy (anonymous):

@gorv help?

OpenStudy (gorv):

cos theta =1/4

OpenStudy (gorv):

cos theta = base / hypotenuse=1/4

OpenStudy (gorv):

base=1 hypotenuse=4 perpendicular=\[\sqrt{4^2-1^2}=\sqrt{16-1}=\sqrt{15}\]

OpenStudy (gorv):

sin theta = perpendicular /hypotenuse = sqrt15/4 as sin theta < 0 sin theta = - sqrt15/4

OpenStudy (gorv):

tan = sin/cos

OpenStudy (gorv):

=-(sqrt15/4)/(1/4)=-sqrt15

OpenStudy (gorv):

first one

OpenStudy (anonymous):

You are literally the best :) can I have your help with just one more?

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

post the new Q

OpenStudy (anonymous):

Okay.